OpenStudy (anonymous):

a question paper has 3 sections . there are 2 question in each section. A candidate has to select two sections and answer one question from each of these selected sections. The number of ways the candidte can answer the question paper is

OpenStudy (anonymous):

@satellite73 @saifoo.khan

OpenStudy (anonymous):

how many ways can the candidate select 2 sections from 3?

OpenStudy (anonymous):

the answer should be 12 @satellite73

OpenStudy (anonymous):

i know the answer, i was asking if you knew how many ways you can pick two sections out of three? we need that number

OpenStudy (anonymous):

i got it as 6

OpenStudy (anonymous):

no, there are only 3 ways to pick two sections out of 3

OpenStudy (anonymous):

suppose the sections are called A, B C then you can pick A B A C B C only three ways

OpenStudy (anonymous):

once you have picked your two sections, you have a choice of two problems in each so total number of ways to do this is \(3\times 2\times 2=12\)

OpenStudy (anonymous):

why 3*2*2???

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

again it is the counting principle figure out how many ways to do each task, then multiply them together

OpenStudy (anonymous):

@satellite73 is it 3*2*1

OpenStudy (anonymous):

3 ways to pick the two sections then in each of the two sections 2 ways to pick the two problems so multiply the number of ways to do each task

OpenStudy (anonymous):

how 2 ways to pick the two problems>>>

OpenStudy (anonymous):

pick the two sections: 3 ways pick a problem from one of the sections chosen: 2 ways pick a problem from the other section chosen: 2 ways \[3\times 2\times 2=12\] it all comes from the "counting principle"

OpenStudy (anonymous):

two ways to pick a problem in a section because you wrote "there are 2 question in each section"

OpenStudy (anonymous):

ok now i get it r u a maths teacher @satellite73

OpenStudy (anonymous):

i am a car

OpenStudy (anonymous):

@satellite73 i am unable to send messages to u....................